Find the value sin2 x on the interval (3pi/2, 2pi) given tan x = -3
Answer by lwsshak3(11628)

Find the value sin2 x on the interval (3pi/2, 2pi) given tan x = -3
tanx=-3/1
hypotenuse of reference right triangle=√(-3)^2+1=√10
sinx=-3/√10
cosx=1/√10
sin(2x)=2sinxcosx=2*(-3/√10)(1/√10)=-6/10=-3/5